# Algebraic non-Hermitian skin effect and unified non-Bloch band theory in arbitrary dimensions

### v1

@@ III.1 Bulk: the non-Bloch layer transfer matrix
We start with a concrete 2D tight-binding model to present our key findings. Importantly, this methodology is universal and applicable to generic 2D systems. The model Hamiltonian in real space is given by
@@
where $c^{\dagger}_{x,y}$ denotes the creation operator of a particle at lattice site $(x,y)$. The Hamiltonian parameters $t_{x},t_{y},t_{xy}$ indicate the hopping strengths along the $x$, $y$, and $x{+}y$ directions, respectively, and $u$ is the on-site potential. These Hamiltonian parameters are generally complex values and thus the Hamiltonian is non-Hermitian and reciprocal. This non-Hermitian Hamiltonian can be implemented in dissipative, reciprocal metamaterials Zhou et al. (2023); Wang et al. (2023a) and quantum systems Zhao et al. (2023). Here, we utilize this reciprocal tight-binding model to demonstrate the effectiveness of our GBZ formula, which addresses the challenges in developing GBZ formulas for reciprocal non-Hermitian systems. We emphasize that our method is applicable to various systems, including non-reciprocal non-Hermitian systems. The lattice size is assumed to be $L_{x}\times L_{y}$. Within the layer-$y$ basis (Fig. 2), the Hamiltonian can be rewritten as
$H=\sum\nolimits_{y}\sum\nolimits_{l=0,\pm 1}\textbf{c}_{y}^{\dagger}\,\textbf{h}_{l}\,\textbf{c}_{y+l},$ (22)
where $\textbf{c}_{y}^{\dagger}$ is a row vector of creation operators at the $y$-th lattice layer, with $L_{x}$ components, expressed as $\textbf{c}_{y}^{\dagger}=(c^{\dagger}_{x=1,y},\dots,c^{\dagger}_{x=L_{x},y})$. Under this layer basis, the Hamiltonian $H$ can be viewed as a band Toeplitz matrix Böttcher and Silbermann (2012) of dimension $L_{y}$ with matrices $\textbf{h}_{l=0,\pm 1}$ as its entries. The matrix elements $\textbf{h}_{l=0,\pm 1}$ themselves are matrices of dimension $L_{x}$ and they represent the couplings between different $y$ layers separated by distance $l$. In total, the Hamiltonian is represented as a matrix of dimension $L_{y}\times L_{x}$. The eigenequation of the Hamiltonian is $H\Psi_{E}=E\Psi_{E}$, and $\Psi_{E}$ is the eigenstate of eigenenergy $E$, represented under the layer basis as $\Psi_{E}=(\psi_{y=1},\dots,\psi_{y=L_{y}})^{T}$. We can cast the bulk eigenequation into a recursion equation along the $y$ direction
@@
where $\mathbb{I}_{L_{x}}$ denotes an identity matrix of dimension $L_{x}$ and $\psi_{y}$ is a column with $L_{x}$ components. Here, $\psi_{y}$ can be further represented as $\psi_{y}=(\Psi_{x=1,y},\dots,\Psi_{x=L_{x},y})^{T}$, where $\Psi_{x,y}$ indicates the wavefunction component at lattice site ($x,y$). The recursion equation can be rewritten in a matrix form
@@
Figure 3:  The GBZ formula and the exact wavefunction constructed from GBZ basis. (a) The blue dots represent the OBC eigenvalues with the system size $L_{x}{=}L_{y}{=}101$. For the chosen generic energy $E_{0}{=}0.819{-}1.108i$ (the red dot), the GFS projections onto the $\beta_{x}$ and $\rho_{y}$ planes are shown in (b) and (c), respectively. (d) The analytical wavefunction for the OBC eigenvalue $E_{0}$ fully agrees with the numerical wavefunction calculated from diagonalizing the Hamiltonian matrix. The Hamiltonian parameters in Eq. (21) are set to be $t_{x}=t_{y}=1,t_{xy}=i/2$, and $u=-i$.
@@
$\begin{split}&f(\beta_{x},\rho_{y},E)=\det{[\mathbb{T}(\beta_{x},E)-\rho_{y}\,\mathbb{I}_{2}]}\\ &=\rho_{y}^{2}+\rho_{y}\,\frac{h_{0}(\beta_{x})-E}{h_{1}(\beta_{x})}+\frac{h_{-1}(\beta_{x})}{h_{1}(\beta_{x})}=0.\end{split}$ (26)
@@
We use the symbol $\rho$ to represent the eigenvalues of the non-Bloch transfer matrix, unless specified otherwise. It is worth noting that the bulk equation Eq. (26) provides constraints on $\beta_{x}$ and $\rho_{y}$, which are fully equivalent to the constraints from the bulk characteristic equation $\det[\mathcal{H}(\beta_{x},\rho_{y})-E]=0$. Here, we introduce the transfer matrix to better solve and understand the wavefunction, while introducing no new constraints. The non-Bloch transfer matrix $\mathbb{T}(\beta_{x},E)$ has two branches of eigenvalues $\rho_{y}$ for each $\beta_{x}$ and energy $E$. These two branches characterize the propagation of the non-Bloch wave $\beta_{x}$ along the $y$ direction. However, the constraints from OBCs have not yet been applied at this stage. We will later demonstrate that not all $\rho_{y}$ branches contribute to the open-boundary eigenstates; instead, only the branches satisfying the GBZ condition are relevant.
III.2 2D GBZ formula and standing-wave construction
Here, we focus on the GBZ formula for the most general 2D case, where $\dim(\sigma_{\text{OBC}})=2$ and $\dim(\text{GBZ})=3$. According to Eq. (13), generating such a three-dimensional GBZ manifold requires a single real GBZ constraint (i.e., a standing wave condition in either one spatial direction). We present the GBZ condition and subsequently derive the GBZ manifold (or equivalently, its submanifold GFS for a specific energy). We demonstrate that in this general scenario, the 2D OBC eigenstates can be constructed using the GBZ basis, represented as a superposition of standing waves along one direction, each coupled with an associated transfer matrix eigenvalue along another direction, as illustrated in Fig. 2. Furthermore, we also discuss in subsection III.5 the other scenarios in two dimensions where the standing wave conditions in both two dimensions are applied and compatible with each other.
@@
We first demonstrate that for a generic 2D lattice Hamiltonian, one can always enforce a standing wave condition along at least one spatial direction. Consider a 2D non-Bloch Hamiltonian $\mathcal{H}(\beta_{x},\beta_{y})$. Without loss of generality, fixing $\beta_{y}$ reduces the 2D Hamiltonian to a 1D subsystem characterized by this fixed $\beta_{y}$. For a given energy $E$, generally, there exist multiple $\beta_{x}$ solutions to the characteristic equation $f(\beta_{x},\beta_{y},E)=0$. These solutions can be organized by their amplitudes as $|\beta_{x,1}(\beta_{y},E)|\leq\dots\leq|\beta_{x,M_{x}+N_{x}}(\beta_{y},E)|$, where $M_{x}$ and $N_{x}$ denote the maximum leftward and rightward hopping ranges, respectively, in the 1D subsystem with fixed $\beta_{y}$. For example, in the model given by Eq. (21), both $M_{x}$ and $N_{x}$ equal 1, indicating that the model only incorporates nearest-neighbor hopping along the $x$ direction for a fixed $\beta_{y}$. The implementation of open-boundary conditions for the 2D system implies that OBCs also apply to each 1D subsystem with a fixed $\beta_{y}$. To reproduce the continuum spectrum for each 1D subsystem under OBCs, the following GBZ condition inherited from the 1D non-Bloch theorem is required:
@@
which imposes a real constraint on ($\beta_{x},\beta_{y},E$). Therefore, at least one standing-wave condition can be imposed in the most generic systems. According to Eq. (13), $\dim{\text{GBZ}}=3$ in this most generic 2D cases.
@@
From the perspective of wavefunctions, in a generic 2D system, the OBC eigenstate can be represented as a superposition of wave components of order-$L$ ($\mathcal{O}(L)$), where $L$ stands for the system length. This is formally expressed as $\Psi_{E}(x,y)=\sum_{(\beta_{x},\beta_{y})}A_{\beta_{x}\beta_{y}}\beta_{x}^{x}\beta_{y}^{y}$. These order-$L$ wave components are typically required to satisfy the order-$L$ degrees of freedom imposed by the open-boundary conditions. It means that, when the system size tends to infinity (in the continuum limit), the solution domain of ($\beta_{x},\beta_{y}$) should be a 1D submanifold embedded in $\mathbb{C}^{2}$. In Hermitian systems, this 1D manifold is the Fermi surface that provides the suitable Bloch wave basis for forming the open-boundary standing waves. However, in 2D non-Hermitian systems, for a fixed energy $E$, with only the constraints from the characteristic equation, $\operatorname{Re}{f}(\beta_{x},\beta_{y},E)=\operatorname{Im}{f}(\beta_{x},\beta_{y},E)=0$, the solution domain of ($\beta_{x},\beta_{y}$) is confined to a 2D surface within the 4D space spanned by $\operatorname{Re}{\beta_{x}},\operatorname{Im}{\beta_{x}},\operatorname{Re}{\beta_{y}},\operatorname{Im}{\beta_{y}}$}. Therefore, an additional constraint—the GBZ condition—is necessary to further reduce this domain to 1D curves, called GFS curves, which enables the proper construction of OBC eigenstates using the non-Bloch basis ($\beta_{x},\beta_{y}$) on the GFS. This GBZ condition is exactly the standing wave condition along either one spatial direction, e.g., the standing wave condition enforced in the $x$ direction as given by Eq. (27). For general 2D systems, we have $\dim{\text{GFS}}=1$, and the OBC spectrum consistently occupies a finite region, i.e., $\dim{\sigma_{\text{OBC}}}=2$. Based on the relationship in Eq. (5), $\dim{\text{GBZ}}=3$ in this most generic 2D cases. Therefore, we can only enforce one standing wave condition along a single spatial direction. The imposition of an additional standing wave condition would decrease the GBZ dimension to $2$ and the GFS dimension to 0, which leads to contradictions. To summarize, in the most generic 2D scenarios, it is permissible and required to enforce just one standing wave condition along either one space dimension. The requirement for two simultaneous standing wave enforcement necessitates more precise adjustments, which we address in subsection III.5.
For a given energy $E_{0}$, the GBZ condition, together with the characteristic equation in Eq. (26), determines the 1D GFS curves in the complex momenta space spanned by $(\beta_{x},\beta_{y})$. As $E_{0}$ sweeps through the entire 2D OBC spectrum, the corresponding GFS constitutes a 3D GBZ manifold. From the GBZ condition in Eq. (27), we see that for a given $\beta_{y}$, the corresponding $\beta_{x}$ solutions can form the standing wave along the $x$ direction.
@@
The GBZ condition in Eq. (27) implies the following general recipe for constructing the OBC eigenstate for a given energy $E$: First, let the non-Bloch wave components $\beta_{x}$ to form standing waves that are satisfying the open boundary conditions in the $x$ direction; then transfer these standing waves along the $y$ direction through the non-Bloch transfer matrix in Eq. (25) and the GBZ condition in Eq. (27). Therefore, each standing wave in the $x$ direction is coupled with the corresponding transfer matrix eigenvalue $\rho_{y}$. Finally, let these standing-wave components superimpose to meet the open boundary conditions in the $y$ direction; the superposition coefficients can be determined and exact open-boundary wavefunction is eventually obtained. Based on this construction recipe, the wavefunction of energy $E$ can be expressed as
$\begin{split}\Psi_{E}(x,y)=\sum\nolimits_{\rho_{y}\in\text{GFS}}A(\rho_{y})\,\rho^{y}_{y}\,\phi_{\rho_{y}}(x),\end{split}$ (28)
@@
where $\phi_{\rho_{y}}(x)$ represents the standing-wave component in the $x$ direction and $A(\rho_{y})$ represents the superposition coefficient. The standing wave component can be further expressed as:
$\phi_{\rho_{y}}(x)=\sum\nolimits_{j=1}^{M_{x}+N_{x}}B_{j}(\rho_{y})\,\beta^{x}_{x,j}(\rho_{y},E),$ (29)
@@
where $B_{j}(\rho_{y})$ represents the superposition coefficients that should satisfy the OBCs in the $x$ direction, and $\beta^{x}_{x,j}(\rho_{y},E)$ satisfies the GBZ condition given in Eq. (27).
@@
We take the Hamiltonian in Eq. (21) as an example, where the Hamiltonian only involves the nearest-neighbor hopping along the $x$ direction, thus $M_{x}=N_{x}=1$. In this model, the GBZ condition under layer-$y$ basis becomes
@@
The open-boundary eigenvalues with system size $L_{x}=L_{y}=101$ are represented by the blue dots in Fig. 3(a), which covers a finite region in the complex energy plane, i.e., $\dim{\sigma_{\text{OBC}}}=2$. We take a generic OBC eigenvalue $E_{0}=0.819-1.108i$, as labeled by the red dot in Fig. 3(a). The corresponding GFS can be obtained according to the GBZ condition in Eq. (30). The projection of GFS onto complex $\beta_{x}$ and $\rho_{y}$ planes are illustrated in Figs. 3(b) and (c), respectively. In our calculation, $2L_{x}$ values of $\rho_{y}$ are obtained as the eigenvalues of the layer-$y$ transfer matrix $\mathbb{T}(E)$ given in Eq. (24). Consequently, these $2L_{x}$ solutions of $\rho_{y}$, represented by blue and red dots in Fig. 3(c), form two arcs in the complex $\rho_{y}$ plane. Here, we emphasize that the 1D GFS curves in the continuum limit can be analytically obtained, similar to solving for the GBZ curves and OBC continuum bands in 1D systems Yang et al. (2020). Specifically, the $\beta_{x}$ and $\rho_{y}$ projection curves [Figs. 3(b) and (c)] of the GFS in a 2D system can be viewed respectively as the GBZ curve and OBC continuum bands of a 1D system Yao and Wang (2018); Yokomizo and Murakami (2019); Zhang et al. (2020); Yang et al. (2020). For a given $\rho_{y}=\rho_{i}$ there are two corresponding $\beta_{x}$ solutions having the same amplitude, i.e., $|\beta_{i,1}(\rho_{i})|=|\beta_{i,2}(\rho_{i})|$, as denoted in Figs. 3(b) and (c). These two $\beta_{x}$ solutions can be used to form a standing wave coupled with $\rho_{i}$. The OBC eigenstate of energy $E_{0}$ can be constructed by these standing-wave components on the GFS. To summarize, for a given energy $E_{0}$, the GBZ condition gives rise to the 1D GFS curve. These non-Bloch waves on the GFS can be used to compose the eigenstate of energy $E_{0}$, as shown in Eq. (28), and the superposition coefficients are determined by the open-boundary conditions in the $y$ direction, as will be shown in the following subsection.
@@
III.3 The exact 2D open-boundary wavefunction from GFS basis

### v2

@@ III.1 Model and non-Bloch layer transfer matrix
We start with a concrete 2D tight-binding model to present our key findings. The model Hamiltonian in real space is given by
@@
where $c^{\dagger}_{x,y}$ denotes the creation operator at lattice site $(x,y)$. The Hamiltonian parameters $t_{x},t_{y},t_{xy}$ indicate the hopping strengths along the $x$, $y$, and x+yx{+}y directions, respectively, and $u$ is the on-site potential. These Hamiltonian parameters are generally complex values and thus the Hamiltonian is non-Hermitian and reciprocal. This non-Hermitian Hamiltonian can be implemented in dissipative, reciprocal acoustic and mechanical metamaterials [84, 85], as well as in ultracold atom systems [90]. Here, we utilize this reciprocal tight-binding model to demonstrate the effectiveness of our GFS formula, though this method is equally applicable to non-reciprocal non-Hermitian systems. A non-reciprocal example is provided in Appendix C.
The reciprocal Hamiltonian in Eq. (21) can be reexpressed in the layer-$y$ basis as follows
$H=\sum\nolimits_{y}\sum\nolimits_{l=0,\pm 1}\textbf{c}_{y}^{\dagger}\,\textbf{h}_{l}\,\textbf{c}_{y+l}$ (22)
with lattice size $L_{x}\times L_{y}$, where $\textbf{c}_{y}^{\dagger}$ is a vector of creation operators at the $y$-th lattice layer, with $L_{x}$ components, expressed as $\textbf{c}_{y}^{\dagger}=(c^{\dagger}_{x=1,y},\dots,c^{\dagger}_{x=L_{x},y})$. Under this layer basis, the Hamiltonian $H$ can be viewed as a band Toeplitz matrix [101] of dimension $L_{y}$ with matrices $\textbf{h}_{l=0,\pm 1}$ as its entries. The matrix elements $\textbf{h}_{l=0,\pm 1}$ themselves are matrices of dimension $L_{x}$ and they represent the couplings between different $y$ layers separated by distance $l$. In total, the Hamiltonian is represented as a matrix of dimension $L_{y}\times L_{x}$. The eigenequation of the Hamiltonian is $H\Psi_{E}=E\Psi_{E}$, and $\Psi_{E}$ is the eigenstate of eigenenergy $E$, represented under the layer basis as $\Psi_{E}=(\psi_{y=1},\dots,\psi_{y=L_{y}})^{T}$. We can cast the bulk eigenequation into a recursion equation along the $y$ direction
@@
where $\mathbb{I}_{L_{x}}$ denotes an identity matrix of dimension $L_{x}$ and $\psi_{y}$ is a column with $L_{x}$ components. Here, $\psi_{y}$ can be further represented as $\psi_{y}=(\psi_{x=1,y},\dots,\psi_{x=L_{x},y})^{T}$, where $\psi_{x,y}$ indicates the wavefunction component at lattice site ($x,y$). The recursion equation can be written in a matrix form
@@
[…]
@@
We use the symbol $\rho$ to represent non-Bloch waves in the transferring direction, and $\beta$ for those in the standing-wave direction. It is worth noting that the bulk equation Eq. (26) provides constraints on $\beta_{x}$ and $\rho_{y}$, which are fully equivalent to the constraints from the bulk characteristic equation $\det[\mathcal{H}(\beta_{x},\rho_{y})-E]=0$. Here, we introduce the transfer matrix to better solve and understand the wavefunction for a finite-size lattice system. The GFS and OBC wavefunctions in the thermodynamic limit can be obtained via an analytic method, which is presented in Appendix A. From Eq. (26), there are two branches of $\rho_{y}$ for given $\beta_{x}$ and energy $E$. They characterize the propagation of the non-Bloch wave $\beta_{x}$ along the $y$ direction in the bulk. However, the standing-wave constraints from OBCs have not yet been applied at this stage. We will in the following section demonstrate that not all $\rho_{y}$ branches contribute to the open-boundary eigenstates; instead, only the branches satisfying the standing-wave condition are relevant.
III.2 GFS formula and standing-wave construction in two dimensions
Here, we focus on generic two-dimensional non-Hermitian systems and demonstrate that $\dim{\sigma_{\text{OBC}}}=2$, $\dim{\text{GBZ}}=3$, and $\dim{\text{GFS}}=1$. According to Eq. (13), constructing the three-dimensional GBZ manifold requires imposing a single constraint, specifically, a standing-wave condition along one spatial direction. We present this standing-wave condition and subsequently derive the resulting GFS at a fixed energy. Based on this, we propose an ansatz for exactly constructing the 2D OBC wavefunction using the GFS basis.
@@ III.2 GFS formula and standing-wave construction in two dimensions
For a generic 2D lattice Hamiltonian, one can enforce the standing wave condition along at least one spatial direction. Consider a 2D non-Bloch Hamiltonian $\mathcal{H}(\beta_{x},\beta_{y})$. Without loss of generality, fixing $\beta_{y}$ reduces the 2D Hamiltonian to a 1D subsystem characterized by this fixed $\beta_{y}$. For a given energy $E$, generally, there exist multiple $\beta_{x}$ solutions to the characteristic equation $f(\beta_{x},\beta_{y},E)=0$. These solutions can be organized by their amplitudes as $|\beta_{x,1}(\beta_{y},E)|\leq\dots\leq|\beta_{x,M_{x}+N_{x}}(\beta_{y},E)|$, where $M_{x}$ and $N_{x}$ denote the maximum leftward and rightward hopping ranges, respectively, in the 1D $\beta_{y}$-subsystem. For example, in the model given by Eq. (21), both $M_{x}$ and $N_{x}$ equal 1, indicating that the model only incorporates nearest-neighbor hopping along the $x$ direction for a fixed $\beta_{y}$. The implementation of open-boundary conditions on the 2D system implies that OBCs each 1D subsystem with fixed $\beta_{y}$ must also satisfy OBCs in the $x$ direction. To reproduce the continuum spectrum for each 1D subsystem under OBCs, the following condition inherited from the 1D GBZ condition is required:
@@
which imposes a real constraint on the complex momentum space of ($\beta_{x},\beta_{y}$) for a given energy $E$. Since this condition involves a fixed energy for the 2D system and determines the GFS curves, we also refer to it as the GFS condition. Therefore, at least one standing-wave condition can be imposed in the most generic systems. According to Eq. (13), $\dim{\text{GBZ}}=3$ in this most generic 2D cases.
@@
From the perspective of wavefunctions, a generic 2D OBC eigenstate can be represented as a superposition of wave components, formally written as $\Psi_{E}(x,y)=\sum_{(\beta_{x},\beta_{y})}A_{\beta_{x}\beta_{y}}\beta_{x}^{x}\beta_{y}^{y}$. The number of such components must match the number of constraints imposed by the open boundary conditions in order to construct the full OBC eigenstate. Thus, the number of components scales with the system length $L$. In the thermodynamic limit, these wave components collectively trace out 1D curves in the 4D space spanned by $(\operatorname{Re}\beta_{x},\operatorname{Im}\beta_{x},\operatorname{Re}\beta_{y},\operatorname{Im}\beta_{y})$. In Hermitian systems, this 1D curve is the Fermi surface that provides the suitable Bloch wave basis for forming the open-boundary eigenstates. However, in 2D non-Hermitian systems, for a fixed energy $E_{0}$, with only the characteristic equation, $\operatorname{Re}{f}(\beta_{x},\beta_{y},E)=\operatorname{Im}{f}(\beta_{x},\beta_{y},E)=0$, the solution domain of ($\beta_{x},\beta_{y}$) is confined to a 2D surface. Therefore, an additional constraint—the standing wave condition—is necessary to further reduce this 2D surface to 1D curves, namely GFS curves. With the GFS basis, the OBC eigenstates can then be constructed exactly. The standing wave condition can be applied along either one spatial direction, e.g., the standing wave condition enforced in the $x$ direction as given by Eq. (27). From dimensionality relation, in general, $\dim{\text{GFS}}=1$ and OBC spectrum occupies a finite region, i.e., $\dim{\sigma_{\text{OBC}}}=2$. Thus, $\dim{\text{GBZ}}=\dim{\text{GFS}}+\dim{\sigma_{\text{OBC}}}=3$. If we impose one more standing wave condition, it would reduce the GBZ dimension to $2$ and the GFS dimension to 0, leading to contradictions that fails to satisfy the constraints imposed by the open-boundary conditions. To summarize, in the most generic 2D setting, it is both permissible and necessary to impose only a single standing-wave condition along one spatial direction. The requirement for two standing wave conditions needs fine-tuning, which we address in subsection III.5.
@@
The GBZ condition in Eq. (27) suggests the following general procedure for constructing the OBC eigenstate at a given energy $E$: First, form standing waves in the $x$-direction by superposing non-Bloch wave components $\beta_{x}$ that satisfy the open-boundary conditions along $x$. Next, propagate these standing waves along the $y$-direction using the non-Bloch transfer matrix defined in Eq. (25), where each x-direction standing wave is associated with a transfer matrix eigenvalue $\rho_{y}$. Finally, superimpose these $y$-propagating components to satisfy the open-boundary conditions in the $y$-direction. The corresponding superposition coefficients are determined by these boundary constraints, yielding the full open-boundary wavefunction. Accordingly, we propose the following ansatz for the 2D OBC wavefunction at energy $E$:
$\begin{split}\Psi_{E}(x,y)=\sum\nolimits_{\rho_{y}\in\text{GFS}}A(\rho_{y})\,\rho^{y}_{y}\,\varphi_{\rho_{y}}(x),\end{split}$ (28)
@@
where $\varphi_{\rho_{y}}(x)$ represents the standing-wave component in the $x$ direction and $A(\rho_{y})$ represents the superposition coefficient. The standing-wave component can be further expressed as:
$\varphi_{\rho_{y}}(x)=\sum\nolimits_{j=1}^{M_{x}+N_{x}}B_{j}(\rho_{y})\,\beta^{x}_{x,j}(\rho_{y},E),$ (29)
@@
where $B_{j}(\rho_{y})$ represents the superposition coefficient that is determined by the OBCs in the $x$ direction, and $\beta^{x}_{x,j}(\rho_{y},E)$ satisfies the GFS condition given by Eq. (27).
Figure 4:  The GFS formula and the exact wavefunction constructed from GFS basis. (a) The blue dots represent the OBC eigenvalues with the system size $L_{x}{=}L_{y}{=}101$. For the chosen generic energy $E_{0}{=}0.819{-}1.108i$ (the red dot), the GFS projections onto the $\beta_{x}$ and $\rho_{y}$ planes are shown in (b) and (c), respectively. (d) The analytical wavefunction for the OBC eigenvalue $E_{0}$ fully agrees with the numerical wavefunction calculated from diagonalizing the Hamiltonian matrix. The Hamiltonian parameters in Eq. (21) are set to be $t_{x}=t_{y}=1,t_{xy}=i/2$, and u=iu=-i.
@@
We take the Hamiltonian in Eq. (21) as an example, where the Hamiltonian only involves the nearest-neighbor hopping along the $x$ direction, thus $M_{x}=N_{x}=1$. In this model, the GFS condition in the layer-$y$ basis becomes
@@
The open-boundary eigenvalues with system size $L_{x}=L_{y}=101$ are represented by the blue dots in Fig. 4(a), which covers a finite region in the complex energy plane, i.e., $\dim{\sigma_{\text{OBC}}}=2$. We take a generic OBC eigenvalue $E_{0}=0.819-1.108i$, as labeled by the red dot in Fig. 4(a). The corresponding GFS can be obtained according to the GFS condition in Eq. (30). The projection of GFS onto complex $\beta_{x}$ and $\rho_{y}$ planes are illustrated in Figs. 4(b) and (c), respectively. We emphasize that by applying the standing-wave condition in Eq. (30) along with the bulk characteristic equation, the GFS in the thermodynamic limit ($L\rightarrow\infty$) can be obtained analytically. The resultant method used to derive the GFS curves shown in Figs. 4(b) and (c) is detailed in Appendix A, and the analytic method for obtaining the GFS in 2D continuum systems is presented in Section VII. The open-boundary conditions in the $x$ direction has been satisfied by the standing-wave condition in the $x$ direction. Therefore, there are OBCs in the $y$ directions leaved. For a finite-size lattice system of size $L_{x}\times L_{y}$, there are $2L_{x}$ lattice points on the two $y$-boundaries, which provide $2L_{x}$ boundary conditions. Meanwhile, the GFS is discretized due to discrete lattice translation symmetry. The discretized GFS should match with the remaining $2L_{x}$ degrees of freedom on the open boundaries in the $y$ direction. Here, we use the layer transfer matrix approach to calculate the GFS in a finite-size lattice system. The layer transfer matrix $\mathbb{T}(E)$, defined in Eq. (24), is a $2L_{x}\times 2L_{x}$ matrix, whose eigenvalues yield the $2L_{x}$ values of $\rho_{y}$. These values represent the discretized GFS curves in the finite size systems. Consequently, these $2L_{x}$ solutions of $\rho_{y}$, represented by blue and red dots in Fig. 4(c), form two arcs in the complex $\rho_{y}$ plane. For a given solution of $\rho_{y}=\rho_{i}$, there are two corresponding $\beta_{x}$ solutions having the same amplitude, i.e., $|\beta_{i,1}(\rho_{i})|=|\beta_{i,2}(\rho_{i})|$, as denoted in Figs. 4(b) and (c). These two $\beta_{x}$ solutions are used to form a standing wave, which coupled with $\rho_{i}$. The OBC eigenstate of energy $E_{0}$ can be constructed by these standing-wave components given by the GFS. To summarize, for a given energy $E_{0}$, the standing-wave condition gives rise to the 1D GFS curve. These non-Bloch waves on the GFS can be used to compose the eigenstate of energy $E_{0}$, as shown in Eq. (28), and the superposition coefficients are determined by the open-boundary conditions in the $y$ direction, as will be shown in the following subsection.
@@
III.3 The 2D wavefunction construction from GFS bases under open boundary conditions